If the statistical odds of getting a disease across the general population stands at 8/100,000. What are the chances against 3 people in a workforce of 130 contracting it?
1
Expert's answer
2013-01-29T09:38:17-0500
The probability for one person of 150 to have this disease is 8/100,000*130.
And the chances of three workers to have a disease are (8/100,000*130)^3 = 0.0104^3 = 0.000001124 or 0.0001124%